Which of the following objects is not experiencing a net force directed north?
(A) An object moving south at a decreasing speed.
(B) An object moving north at an increasing speed.
(C) An object instantaneously at rest that then begins to move north.
(D) An object moving north at a constant speed.

Grade:upto college level

1 Answers

Navjyot Kalra
8 years ago
The correct option is (D) an object moving north at a constant speed.
In accordance with Newton’s Second Law of Motion, force (F) is equal to the rate of change of momentum (p)
F = dp/dt …… (1)
Momentum of an object (p) is equal to the mass of the particle (m) time’s velocity of the object (v).
p = mv …… (2)
Putting the value of equation (2) in equation (1), we get,
F = dp/dt
= d(mv)/dt
= m dv/dt …… (3)
If an object moves with a constant speed (v = constant), then the term dv/dt in equation (3) will be equal to zero.
So, dv/dt = 0 …… (4)
Putting the value of equation (4) in equation (3), force will be,
F = m dv/dt
= (m) (0)
= 0 …… (5)
Therefore an object moving north at a constant speed is not experiencing a net force directed north.
From the above observation we conclude that, option (D) is correct.